Flint Company sells tablet PCs combined with Internet service, which permits the tablet to connect to the Internet anywhere and set up a Wi-Fi hot spot. It offers two bundles with the following terms.

1. Flint Bundle A sells a tablet with 3 years of Internet service. The price for the tablet and a 3-year Internet connection service contract is $512. The standalone selling price of the tablet is $246 (the cost to Flint Company is $174). Flint Company sells the Internet access service independently for an upfront payment of $312. On January 2, 2020, Flint Company signed 100 contracts, receiving a total of $51,200 in cash.
2.

Flint Bundle B includes the tablet and Internet service plus a service plan for the tablet PC (for any repairs or upgrades to the tablet or the Internet connections) during the 3-year contract period. That product bundle sells for $619. Flint Company provides the 3-year tablet service plan as a separate product with a standalone selling price of $160. Flint Company signed 200 contracts for Flint Bundle B on July 1, 2020, receiving a total of $123,800 in cash.

Repeat the requirements for part (a), assuming that Flint Company has no reliable data with which to estimate the standalone selling price for the Internet service
Answer #1

All amounts are in $

There are two components (two performance obligations) in the first bundle

There are three components (three performance obligations) in second bundle

Calculating each performance price

Total of Bundle 1 (seperately) = 246+312 = 558

Total of Bundle 2 (separately) = 246+312+160 = 718

In Bundle 1 :

Tablet = 512 x (246/558) = 225.72

Internet Service = 512 x (312/558) = 286.28

In Bundle 2 :

Tablet = 619 x (246/718) = 212.08

Internet Service = 619 x (312/718) = 268.98

Maintenance service = 619 x (160/718) = 137.94

Journal entries Part on Jan 2, 2020 Cash $51,200 Sales revenue $ 22,572 (225727100) Uhlemnes sens, revenue $281628 (286.28x10price is If there reliable data, the Internet service $512 - $246 $266 Conly in part 1) On Jan 2, 2020 Cash $51,200 Sales rev

Note :

1. For 2nd part, the revenue is recognised for only 6 months in first year (because sold in July) and the revenue is recognised for 12 months in next two years. Later again it is recognised for remaining 6 months in 4th year

2. The Cost of goods sold entry is accounted on the same day of sale (Accounts maintained as per perpetual inventory will be accounting these entries. As no information is available on the type of acounts maintained, these entries are accounted separately for your better understanding)

3.

Accounts used in this problem are

Sales Revenue

Service Revenue

Sales Revenue (Internet)

Sales Revenue (maintenance)

Unrecognised Service Revenue

Unrecognised Service Revenue (Internet)

Unrecognised Service Revenue (maintenance)

Cash

Cost of goods sold

Inventory